### v1

# 2Zhongguancun Laboratory, Beijing, China

Abstract
LLM Ensemble—which involves the comprehensive use of multiple large language models (LLMs), each aimed at handling user queries during downstream inference, to benefit from their individual strengths—has gained substantial attention recently. The widespread availability of LLMs, coupled with their varying strengths and out-of-the-box usability, has profoundly advanced the field of LLM Ensemble. This paper presents the first systematic review of recent developments in LLM Ensemble. First, we introduce our taxonomy of LLM Ensemble and discuss several related research problems. Then, we provide a more in-depth classification of the methods under the broad categories of “ensemble-before-inference, ensemble-during-inference, ensemble-after-inference”, and review all relevant methods. Finally, we introduce related benchmarks and applications, summarize existing studies, and suggest several future research directions. A curated list of papers on LLM Ensemble is available at https://github.com/junchenzhi/Awesome-LLM-Ensemble.
{NoHyper}footnotetext: *Corresponding authors.
1 Introduction
In recent years, the landscape of artificial intelligence has been dramatically reshaped by the development of Large Language Models (LLMs), including Gemini Team et al. (2023), GPT-4 Achiam et al. (2023), Llama Touvron et al. (2023), and the recently introduced DeepSeek Liu et al. (2024a). The success of these LLMs continues to fuel widespread research enthusiasm, with a remarkable total of over 182,000 large language models now accessible in the Hugging Face library.  [^1] https://huggingface.co/models.
Behind this research enthusiasm, however, we can identify two main aspects: 1) The performance concerns: The direct out-of-the-box capability of LLMs (from zero-shot inference) and their indirect out-of-the-box capability (from in-context-learning few-shot inference) still raise performance worries, including accuracy, hallucinations, and misalignment with human intent, among others; 2) The varying strengths and weaknesses of LLMs, each with different inference costs: Due to differences in architecture, parameter size, tokenization, dictionary, training data, and methodology, these LLMs exhibit substantial variability and their responses can differ significantly. With the above two aspects in mind and drawing on the spirit of Ensemble Learning, it is natural to consider that, for each task query, rather than persistently relying on a single LLM based on public rankings or other criteria, it might be more advantageous to simultaneously consider multiple LLM candidates (usable out-of-the-box) and harness their distinct strengths. This is exactly what the recently emerging field of LLM Ensemble explores.
(a) Ensemble before inference.
(b) Ensemble during inference.
(c) Ensemble after inference.
Figure 1: Illustration of the LLM Ensemble taxonomy. (Note that for (b) ensemble-during-inference, there is also a (b3) process-level ensemble approach not depicted in the figure, due to layout considerations and the status that this approach is instantiated by a single method.)
Existing LLM Ensemble methods can be broadly categorized into three types, depending on the sequence of LLM inference and ensemble: 1) Ensemble-before-inference approach, utilizes the given query information while considering the diverse characteristics of all LLM candidates to route an appropriate model for inference (this approach is similar to the hard voting strategy in Ensemble Learning); 2) Ensemble-during-inference approach, aggregates incomplete responses (e.g., token-level information) from multiple LLMs during the decoding process and feeds the combined result back into all the models; 3) Ensemble-after-inference approach, performs the ensemble after full responses (instead of fragments) have been generated by all models or a subset of them. Despite the emergence of numerous methods derived from these broad paradigms recently, there is still no formal survey that offers a comprehensive review of the core ideas and related research in the rapidly evolving field of LLM Ensemble.
We present the first comprehensive survey on LLM Ensemble, introducing recent advances and focusing on taxonomy, related problems, methods, benchmarks, applications, and future directions. We hope that this survey will provide a thorough review for researchers and inspire further exploration.
2 LLM Ensemble Taxonomy and Related Problems
2.1 LLM Ensemble Taxonomy
This section formally introduces our LLM Ensemble taxonomy, illustrated by the schematic shown in Figure 1 and the detailed taxonomy in Figure 2. As mentioned in Section 1, the following three broad categories of LLM Ensemble exist.
(a) Ensemble before inference. In essence, this approach employs a routing algorithm prior to LLM inference to allocate a specific query to the most suitable model, allowing the selected model that is specialized for the query and typically more cost-efficient inference to perform the task. As illustrated in Figure 1 and Figure 2, existing methods can be classified into two categories, depending on whether the router necessitates the use of pre-customized data for pre-training: (a1) pretrained router and (a2) non-pretrained router.
(b) Ensemble during inference. As the most granular form of ensemble among the three broad categories, this type of approach encompasses: (a1) token-level ensemble methods, integrate the token-level outputs of multiple models at the finest granularity of decoding; (a2) span-level ensemble methods, perform ensemble at the level of a sequence fragment (e.g., a span of four words); (a3) process-level ensemble methods, select the optimal reasoning process step-by-step within the reasoning chain for a given complex reasoning task. Note that for these ensemble-during-inference methods, the aggregated text segments will be concatenated with the preceding text and fed back into the models.
(c) Ensemble after inference. These methods can be classified into two categories: (c1) Non-cascade methods, perform ensemble by integrating multiple complete responses contributed from all LLM candidates; (c2) Cascade methods, consider both performance and inference costs, progressively performing inference through a chain of LLM candidates ranked primarily by model size to identify the most suitable response and terminate the cascade process.
2.2 Related Problems
Here we briefly introduce the closely related problems.
LLM Merging, also known as LLM Fusion Yang et al. (2024), integrates parameters from various LLMs to construct a universal model without requiring original training data or extensive computation. It is highly relevant to LLM Ensemble, as both promote knowledge fusion and transfer.
LLM Collaboration leverages the distinct strengths of each model to approach tasks with increased flexibility Du et al. (2023); Lu et al. (2024a). Unlike LLM Ensemble where models are employed with equal status to directly faced to user queries, the collaboration approach assigns distinct roles to each LLM, exchanging response information to enhance task resolution.
Weak Supervision Zhang et al. (2021); Chen et al. (2023c), also known as Learning from Crowds Chen et al. (2021, 2022), uses weak labels to perform information aggregation Chen et al. (2023b) (corresponding non-cascade ensemble after inference in LLM Ensemble) or directly train a classifier Zhang et al. (2021), yet most methods focus on classification, rather than general generation tasks.
3 Methodology
In this section, following the taxonomy in Section 2.1, we systematically review the three types of methods—ensemble before inference, ensemble during inference and ensemble after inference—in Sections 3.1,  3.2, and  3.3, respectively.
{forest}
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw, rounded corners,align=left, minimum width=2.5em, inner xsep=4pt, inner ysep=1pt, , where level=1text width=5em,fill=blue!10, where level=2text width=5em,font=,fill=pink!30, where level=3font=,yshift=0.26pt,fill=yellow!20, [LLM
Ensemble
methods,fill=green!20 [(a) Ensemble-before-
inferece (Sec. 3.1),text width=7.6em [(a1) Pretrained router,text width=7.8em [Classification-based: Shnitzer et al. (2023) / Srivatsa et al. (2024) / Ong et al. (2024) / Hybrid-LLM Ding et al. (2024) /
SelectLLM Maurya et al. (2024) / Routoo Mohammadshahi et al. (2024) ] [Reward-based: Zooter Lu et al. (2024b) / Bench-CoE Wang et al. (2024) / Li (2025) / MetaLLM Nguyen et al. (2024) /
TO-Router Stripelis et al. (2024) ] [Assignment-based: FORC Šakota et al. (2024) / HomoRouter Mu et al. (2024) ] ] [(a2) Non-pretrained router,text width=9.8em [Eagle Zhao et al. (2024) / PickLLM Sikeridis et al. (2024) / Blending Lu et al. (2024c) ] ] ] [(b) Ensemble-during-
inference (Sec. 3.2),text width=7.6em [(b1) Token-level ensemble,text width=9.6em [Aggregation-based: 1) Vocabulary alignment: GaC Yu et al. (2024) / DeePEn Huang et al. (2024) /
EVA Xu et al. (2024) / 2) Others: UniTe Yao et al. (2024) / PackLLM Mavromatis et al. (2024) / CDS Jin et al. (2024) ] [Selection-based: CITER Zheng et al. (2025) ] ] [(b2) Span-level ensemble, text width=9.6em [Cool-Fusion Liu et al. (2024b) / SweetSpan Xu et al. (2025) / SpecFuse Lv et al. (2024b) ] ] [(b3) Process-level ensemble, text width=10.1em [LE-MCTS Park et al. (2024) ] ] ] [(c) Ensemble-after-
inference (Sec. 3.3),text width=7.6em [(c1) Non-cascade,text width=6.4em [Selection-based: Agent-Forest Li et al. (2024) / Smoothie Guha et al. (2024) / MoRE Si et al. (2023)] [Selection-then-regeneration: LLM-Blender Jiang et al. (2023) / URG Lv et al. (2024a) / LLM-TOPLA Tekin et al. (2024) ] ] [(c2) Cascade, text width=5.4em [Unsupervised: EcoAssistant Zhang et al. (2023) / Yue et al. (2024) / Model Cascading Varshney and Baral (2022) /
neural caching Ramírez et al. (2023) / Cascade Routing Dekoninck et al. (2024) ] [Supervised:  Jitkrittum et al. (2024) / Gupta et al. (2024) / FrugalGPT Chen et al. (2023a) / AutoMix Aggarwal et al. (2023) /
DER Hu et al. (2024a) / ] ] ] ]
Figure 2: Taxonomy of LLM Ensemble methods.
3.1 Ensemble Before Inference
As mentioned in Section 2.1, two categories of ensemble-before-inference approaches exist: pretrained router and non-pretrained router. (a1) Pretrained router approach can be further classified into: classification-based methods, reward-based methods and assignment-based methods. For these methods, to effectively learn and deploy a pretrained router model, a series of critical phases need to be executed, from data preparation to router model training, evaluation and router model deployment/serving. (a2) The non-pretrained router approach designs some selection strategies to build the router model. Below we present a review of these methods and provide a comprehensive summary in Table 1.
3.1.1 3.1.1 (a1) Pretrained Router
Classification-based methods.
Shnitzer et al. Shnitzer et al. (2023) are the first to present the pretrained router model for selecting the most suitable LLM for a given query. They transform the model selection problem into multiple binary classification problems, where each aims to predict whether a specific LLM will correctly respond to an input query. Theoretically, Srivatsa et al. Srivatsa et al. (2024) adopt classifier-based and clustering-based routing methods, introducing theoretical upper bounds to evaluate the models. Similarly, Ong et al. Ong et al. (2024) formulate the LLM routing problem to explore the trade-off between cost and response quality. Their router training framework, based on human preference data and augmentation techniques, demonstrates over 2$\times$ cost savings on widely used benchmarks. Hybrid-LLM Ding et al. (2024) shares similarities with the framework of Ong et al. Ong et al. (2024) but differs in three key ways: it uses synthetic preference labels derived via BARTScore, relies on a single BERT-based router architecture, and limits evaluation to in-domain generalization. SelectLLM Maurya et al. (2024) uses a multi-label classifier to model the capabilities of different LLMs and selects LLMs through confidence-based policies. Routoo Mohammadshahi et al. (2024) decomposes the LLM router problem into two core components: classification-based performance prediction and cost-aware selection. Through their collaboration, these components enable Routoo to route queries to the most appropriate LLM, balancing cost and performance.
Methods Un-/Supervised Goals Loss functions Tasks Generalization Code
(a1) Shnitzer et al. (2023) Supervised Performance Binary CE loss Generation -
Srivatsa et al. (2024) Un-/Supervised Performance Class-balanced CE loss/Clustering Reasoning [Link]
Ong et al. (2024) Supervised Performance and cost Binary CE loss Generation [Link]
Hybrid-LLM Ding et al. (2024) Supervised Performance and cost Binary CE loss Generation [Link]
SelectLLM Maurya et al. (2024) Supervised Performance, cost Class-balanced CE loss Understanding/Reasoning -
Routoo Mohammadshahi et al. (2024) Supervised Performance and cost CE loss Classification -
Zooter Lu et al. (2024b) Supervised Performance KL divergence Generation -
Bench-CoE Wang et al. (2024) Supervised Performance CE loss Generation [Link]
Li (2025) Supervised Performance and cost Negative log-likelihood loss Generation -
MetaLLM Nguyen et al. (2024) Supervised Performance Mean square error Classification [Link]
TO-Router Stripelis et al. (2024) Supervised Perf., cost and query time KL divergence Generation -
FORC Šakota et al. (2024) Supervised Performance and cost Avg. of multi-metric integration Generation [Link]
HomoRouter Mu et al. (2024) Supervised Performance and cost Mean square error Generation -
(a2) Eagle Zhao et al. (2024) Unsupervised Performance and cost - Generation -
PickLLM Sikeridis et al. (2024) Supervised Perf., cost and query time - Generation -
Blending Lu et al. (2024c) Unsupervised Performance and cost - Conversation -
Table 1: Summary of ensemble-before-inference methods.
Reward-based methods.
Zooter Lu et al. (2024b) is an efficient reward-guided routing method for integrating existing LLMs. It uses the scores of reward models as supervision signals to train the routing function through a reward distillation mechanism, achieving efficient routing of queries to specialized LLMs. Unlike previous fine-grained routing methods, Bench-CoE Wang et al. (2024) takes into account both query-level and subject-level routing strategies. To address the performance-cost dilemma, Li et al. Li (2025) formulate the LLM selection process as a multi-armed bandit problem, optimizing the trade-off between performance and cost. MetaLLM Nguyen et al. (2024) optimizes a scalarized reward and operates on a fixed set of LLMs, limiting the learned policy to specific user preferences and a predefined set of models. TO-Router system Stripelis et al. (2024) effectively balances query execution time, cost, and model performance through dynamic routing, integrating multiple LLM experts into a unified query interface.
Assignment-based methods.
FORC Šakota et al. (2024) provides a variety of assignment strategies, including both cost-insensitive strategies and cost-sensitive strategies. Users can choose the appropriate strategy according to different needs and constraints to achieve the best balance between cost and performance. In the retrieval-augmented generation (RAG) scenario, given a set of homogeneous search tools and user queries, HomoRouter Mu et al. (2024) is trained to predict the score of the LLM calling each tool to process the query, and then the query is assigned to the best tool according to different allocation strategies.
3.1.2 3.1.2 (a2) Non-Pretrained Router
Unlike learning-based routers, Eagle Zhao et al. (2024) is a training-free router, which employs the ELO ranking to update the scores of LLMs based on pairwise comparison results. For each comparison, the observed scores and the expected scores are calculated to determine the score changes of the LLMs, thereby evaluating the relative skill levels of the LLMs and providing the highest quality responses to users within the given budget. PickLLM Sikeridis et al. (2024) adopts reinforcement learning to dynamically route queries to the most appropriate LLMs. By designing a weighted reward function, it comprehensively considers multiple factors such as cost, query time, and response accuracy. Blending Lu et al. (2024c) introduces a simple yet effective conversational AI method, which randomly selects different responses of multiple chat AIs, breaking the traditional model of relying on a single chat model. It effectively combines the advantages of multiple small models and significantly improves the performance of chat AI while maintaining a low inference cost.
Methods Granularities Main Strategies Unsup. # Models Other Traits§ Efficiency-Aware Tasks Code
(b1) GaC Yu et al. (2024) Token Averaging agg.,Vocabulary align. $\geq 2$ Union dictionary Generation [Link]
DeePEn Huang et al. (2024) Token Aggregation,Vocabulary align. $\geq 2$ Relative representation Generation [Link]
EVA Xu et al. (2024) Token Averaging agg.,Vocabulary align. $\geq 2$ Relative representation Generation [Link]
UniTe Yao et al. (2024) Token Averaging agg.,TOP-K union $\geq 2$ Union dictionary Generation -
PackLLM Mavromatis et al. (2024) Token Weighted avg. agg.,Perplexity $\geq 2$ Greedy optimization Generation [Link]
CDS Jin et al. (2024) Token Weighted-averaging agg. $\geq 2$ Critical token classifier Generation -
CITER Zheng et al. (2025) Token Token-Level routing (selection) $=2$ Reinforcement learning Generation [Link]
(b2) Cool-Fusion Liu et al. (2024b) Span Generation-assessment-selection $\geq 2$ Perplexity, common words Generation -
SweetSpan Xu et al. (2025) Span Generation-assessment-selection $\geq 2$ Perplexity, fixed-length Generation -
SpecFuse Lv et al. (2024b) Span Generation-assessment-selection $\geq 2$ Perplexity, fixed-length Generation -
(b3) LE-MCTS Park et al. (2024) Process Reasoning process selection $\geq 2$ Monte Carlo Tree Search Reasoning -
  • 1
    §$\S$: “Other Traits§” denotes unique and relatively important capabilities, modules or characteristics that are noteworthy.
  • 2
[…]
Table 2: Summary of ensemble-during-inference methods.
3.2 Ensemble During Inference
We present a comprehensive review of token-level, span-level, and process-level ensemble methods in the following subsections and offer a summary analysis in Table 2.
3.2.1 3.2.1 (b1) Token-Level Ensemble
For token-level ensemble-during-inference methods, at each decoding time-step, 1) aggregation approach Yu et al. (2024); Huang et al. (2024); Xu et al. (2024); Yao et al. (2024); Mavromatis et al. (2024); Jin et al. (2024) produces the final token distribution for generation by (weighted) averaging token-level probability distributions from different models, whereas 2) selection approach Park et al. (2024) opts to directly adopt the output token from a selected single model.
Aggregation-based methods.
When attempting token-level aggregation, one encounters a dilemma—the vocabulary discrepancies across different LLMs, which produce varying embedding lengths, impede the fusion and averaging of multiple probability distributions, creating a substantial obstacle during ensemble.
Thus, several methods Yu et al. (2024); Huang et al. (2024); Xu et al. (2024) focus on addressing the vocabulary alignment issue. (i) GAC Yu et al. (2024) is the most straightforward method, which constructs a new “union dictionary” by combining the vocabularies of multiple models to include all tokens from each dictionary, and subsequently projects the distribution information derived from each model onto this new merged dictionary for averaging aggregation. (ii) Further, DeePEn Huang et al. (2024) and EVA Xu et al. (2024) project the output distributions of multiple models at the current time step into a shared relative/pivot space, followed by either an averaging aggregation Huang et al. (2024) or a weighted aggregation Huang et al. (2024); Xu et al. (2024) to derive an unified final distribution, where the mapping process is grounded in the relative representations theory Moschella et al. (2022). Under this same idea, one key distinction between the two methods lies in their approach to constructing the dictionary for the relative/pivot space.
Furthermore, there are several alternative methods that do not prioritize dictionary alignment as their core objective: UniTe Yao et al. (2024) and PackLLM Mavromatis et al. (2024) directly adopt existing alignment techniques, and CDS Jin et al. (2024) employs homogeneous models in their context without the need for alignment. Specifically: (i) Yao et al. Yao et al. (2024) propose UniTe, which focuses solely on the TOP-K portion of each model’s output distribution and employs the aforementioned union dictionary-based alignment strategy Yu et al. (2024) for performing averaging aggregation, avoiding the need for full vocabulary alignment and reducing computational overhead. (ii) Mavromatis et al. Mavromatis et al. (2024) present a weighted averaging approach, where the weights were determined by the perplexity of each model at that specific time step, serving as an indicator of whether the current input aligns with the respective model’s expertise. Here we present the formal definition of perplexity (considering that it is a crucial concept also employed in several other related methods Liu et al. (2024b); Xu et al. (2025)). Given a tokenized input prompt or sequence $S=(x_{0},x_{1},\ldots,x_{t})$,
$\operatorname{PPL}_{k}(S)=\exp\{-\frac{1}{t}\sum_{i}^{t}\log p_{k}(x_{i}\mid x_{<i})\},$ (1)
where the perplexity value is determined by the exponentiated average negative log-likelihood of $S$, and $\log p_{k}(x_{i}\mid x_{<i})$ denotes the log-likelihood of $i$-token given the preceding tokens $x_{<i}$ according to model $M_{k}$. Further, a relatively complex variant of this approach is derived from a perplexity-based greedy optimization problem. (iii) CDS Jin et al. (2024) ensembles the outputs of multiple homogeneous models (derived from a single ancestor model after pre-training or fine-tuning/alignment) at key decoding points (determined by a critical token classifier), enabling the simultaneous utilization of their complementary advantages to enhance the factuality of the generated content.
Selection-based methods.
As a separate class of approaches, Zheng et al. Zheng et al. (2025) train a router driven by reinforcement learning to determine, at each token-level step, whether the output from a small model or a potentially triggered subsequently larger model should be selected, offering an efficiency-aware solution.
3.2.2 3.2.2 (b2) Span-Level Ensemble
Existing span-level ensemble methods  Liu et al. (2024b); Xu et al. (2025); Lv et al. (2024b) all adhere to a “generation-assessment-selection” pipeline. These methods first employ multiple large models to generate word-containing fragments, then utilize the concept of perplexity, as mentioned before (Equation 1), to make each model score all generated model responses, ultimately selecting the fragment with the highest cumulative score. The key distinction lies in that Gool-Fusion Liu et al. (2024b) have each source LLM individually generate text segments until the word boundary of each segment is common to all LLMs, whereas SweetSpan Xu et al. (2025) and SpaceFuse Lv et al. (2024b) adopt a more straightforward approach by setting the span as a fragment with a fixed word count. Furthermore, SpaceFuse Lv et al. (2024b) introduces a model exit mechanism during each task execution phase to optimize computational resource usage.
3.2.3 3.2.3 (b3) Process-Level Ensemble
Confronted with complex step-by-step reasoning tasks, Park et al. Park et al. (2024) use a trained Monte Carlo Tree Search strategy at each reasoning step to select the output with the highest reward value from multiple model reasoning outputs, thereby identifying the most accurate reasoning chain.
Methods Un-/Supervised Main Strategies # Models Other Traits§ Efficiency-aware Tasks Code
(c1) Agent-Forest Li et al. (2024) Unsupervised Similarity-based selection $>2$ Scaling property Cls./Generation [Link]
Smoothie Guha et al. (2024) Unsupervised Similarity-based selection $>2$ Graphical model Generation [Link]
MoRE Si et al. (2023) Supervised Supervised Sim.-based sel. $>2$ Feature engineering Reasoning [Link]
LLM-Blender Jiang et al. (2023) Supervised Selection-then-regeneration $>2$ Pairwise ranking Generation [Link]
LLM-TOPLA Tekin et al. (2024) Supervised Selection-then-regeneration $>2$ Maximising diversity Generation [Link]
URG Lv et al. (2024a) Supervised Selection-then-regeneration $>2$ Unified processes Cls./Generation -
(c2) EcoAssistant Zhang et al. (2023) Unsupervised User judgment $\geq 2$ Prompt engineering Code generation [Link]
Yue et al. (2024) Unsupervised Answer consistency $\geq 2$ Sampling and checking Reasoning [Link]
Model Cascading Varshney and Baral (2022) Unsupervised Class uncertainty $\geq 2$ Cascade-pioneer Classification -
neural caching Ramírez et al. (2023) Unsupervised Class unc./Ans. cons. $=2$ Distillation Classification [Link]
Cascade Routing Dekoninck et al. (2024) Unsupervised Class uncertainty $>2$ Routing-equipped Classification [Link]
Jitkrittum et al. (2024) Supervised (Upgraded) Class unc. $\geq 2$ Post-hoc deferral Classification -
FrugalGPT Chen et al. (2023a) Supervised Score function $\geq 2$ Prompt engineering, etc. Query answering -
Gupta et al. (2024) Supervised Score function $=2$ Quantile features Generation -
AutoMix Aggarwal et al. (2023) Supervised MDP $\geq 2$ Routing-equipped, etc. Reasoning -
DER Hu et al. (2024a) / Supervised MDP,Ans. Cons. $\geq 2$ Routing-equipped Reasoning -
  • 1
    §$\S$: “Other Traits§” denotes unique and relatively important capabilities, modules or characteristics that are noteworthy. Beyond the introduction already provided in the main text, Zhang et al. Zhang et al. (2023)
    employ some prompt engineering, and Yue et al. Yue et al. (2024) use the customized sampling and checking techniques.
  • 2
[…]
Table 3: Summary of ensemble-after-inference methods.
3.3 Ensemble After Inference
We review the non-cascade and cascade methods in the following subsections, and provide a summary analysis in Table 3.
3.3.1 3.3.1 (c1) Non-Cascade
As shown in Figure 2 and Table 3, existing non-cascade methods fall into two categories: 1) selection-based, which is dedicated to selecting a single response from multiple candidates Li et al. (2024); Guha et al. (2024); Si et al. (2023), and 2) selection-then-regeneration, which involves initially selecting a subset of candidate responses and subsequently feeding this refined subset into a generative model for regeneration to obtain the final output Jiang et al. (2023); Tekin et al. (2024); Lv et al. (2024a).
Selection-based methods.
First, as unsupervised methods, Agent-Forest Li et al. (2024) and Smoothie Guha et al. (2024) essentially leverage the similarity between model responses for selection and adhere to a majority voting (MV) principle: selecting the response that exhibits the highest degree of similarity to the others as the final answer. Specifically, (i) Li et al. Li et al. (2024) obtain multiple responses by repeatedly querying a single model (equivalent to utilizing outputs from multiple homogeneous models) and employ the MV ensemble, discovering a scaling property associated with ensembling more responses; they use BLEU scores and occurrence frequencies to quantify the similarities between responses for generation and classification tasks, respectively. (ii) Guha et al.  Guha et al. (2024) input multiple $<$query, response$>$ pairs into SentenceBERT Reimers (2019), leveraging the Euclidean distance between the resulting low-dimensional encoded vectors as a measure of similarity; and their method is theoretically supported by a constructed graphical model.
Further, as a supervised method, MoRE Si et al. (2023) utilizes some training data to train a random forest classifier for selecting among multiple responses, using the response similarity among LLMs as one of the constructed features.
Selection-then-regeneration methods.
These methods stem from the pioneering work LLM-Blender, introduced by Jiang et al. Jiang et al. (2023). (i) LLM-Blender uses some training data to train the “PairRanker” selection module in the first phase, serving to select a subset from multiple responses, and to train the “GenFuser” generator in the subsequent phase for synthesizing the final response. (ii) Building upon LLM-Blender, the improvement made by LLM-TOPLA Tekin et al. (2024) mainly lies in optimizing the selection process in the first step by “maximizing the diversity among multiple response candidates”. (iii) In addition, URG Lv et al. (2024a) proposes an end-to-end framework that integrates selection and regeneration, aiming to avoid the time consumption and error propagation caused by the original two separate processes.
3.3.2 3.3.2 (c2) Cascade
Methodologically, all cascade methods revolve around the deferral rule/decision maker: determining whether to adopt the output of the current model at hand or to invoke a subsequent, more powerful model. As shown in Figure 2 and Table 3, this line of methods can be classified into two categories: unsupervised methods and supervised methods.
Unsupervised methods.
As shown in Table 3, these methods can be further categorized into three types according to the pivotal “main strategy”: user judgment Zhang et al. (2023), answer consistency Yue et al. (2024); Ramírez et al. (2023), and class uncertainty Varshney and Baral (2022); Ramírez et al. (2023); Dekoninck et al. (2024).
As one most straightforward method, Zhang et al. Zhang et al. (2023) propose a cascade method, EcoAssistant, that relies on user judgment rather than machine algorithms to determine whether the obtained result are satisfactory and whether the cascading inference should be terminated.
Further, Yue et al. Yue et al. (2024) introduce a method based on answer consistency—more precisely, on the observation of whether the inference results generated from multiple identical/diverse prompts exhibit sufficient uniformity—to solve the cascade problem for in-context-learning reasoning tasks. The underlying intuition is that, if a weaker LLM produces highly consistent answers for a given task, it demonstrates strong confidence in solving the problem, and its most consistent answer is likely correct; therefore, invoking a stronger LLM is unnecessary. In addition, the strategies provided by the method “neural caching” Ramírez et al. (2023) also contain the answer consistency.
Last but certainly not least, a critically important category of approaches is grounded in class uncertainty——an approach that also manifests in the following “supervised methods”. Class uncertainty, sometimes also referred to as confidence Varshney and Baral (2022); Ramírez et al. (2023); Jitkrittum et al. (2024), can be implemented through various variants, including Maximum Softmax Probability Varshney and Baral (2022), Distance To Uniform Distribution Varshney and Baral (2022), Margin Sampling Ramírez et al. (2023), and Prediction Entropy Ramírez et al. (2023). However, at its core, it evaluates whether the probability of the dominant class inferred by the current model exceeds a sufficient threshold. When this probability is sufficiently high—signifying adequate confidence—the cascading process is concluded; otherwise, it proceeds. Specifically, (i) “Model Cascading” Varshney and Baral (2022) pioneers this approach. (ii) Ramírez et al. Ramírez et al. (2023) propose the “neural caching” embedded the class uncertainty strategy. Its cascading process involves only two models, and during the task execution phase, the weaker model continuously learns from the distillation outputs of the stronger model to enhance its capabilities for better handling subsequent tasks. (iii) Dekoninck et al.  Dekoninck et al. (2024) apply the routing approach in the cascading process and theoretically prove that, under certain conditions, their method essentially adopts the class uncertainty strategy.
Supervised methods.
Unlike the aforementioned unsupervised cascade methods, supervised cascade methods use some training data to train cascade-related modules, including post-hoc deferral based on upgraded class-uncertainty strategy Jitkrittum et al. (2024), scoring functions Chen et al. (2023a); Gupta et al. (2024), and Markov Decision Processes (MDPs) Aggarwal et al. (2023); Hu et al. (2024a).
Approaches Ensemble Strategies Ensemble Granularities Ensemble Goals
[…]
  • 1
[…]
Table 4: Summary analysis of the key attributes of LLM Ensemble approaches.
Specifically, (i) Jitkrittum et al. Jitkrittum et al. (2024) use the training data to learn a cascade method called post-hoc deferral; and the core idea is that, in contrast to aforementioned class-uncertainty cascade methods Varshney and Baral (2022); Ramírez et al. (2023); Dekoninck et al. (2024) which solely focus on the class uncertainty of the current model’s output, it also takes into account the estimated uncertainty of the next, stronger model’s output. (ii) Further, Chen et al. Chen et al. (2023a) and Gupta et al. Gupta et al. (2024) use the training data to learn a scoring function that produces a confidence score regarding the current model’s output, enabling cascading judgments. The key difference lies in that Chen et al. Chen et al. (2023a) utilize input features contributed by $<$query, response$>$ pairs during the scoring function learning process; whereas Gupta et al. Gupta et al. (2024), aiming to better integrate the class-uncertainty idea in general generation tasks, employ “Quantile” features constructed from token-level class uncertainty information within the response in the binary classification-based scoring function learning. (iii) Both Aggarwal et al. Aggarwal et al. (2023) and Hu et al. Hu et al. (2024a) adopt the aforementioned idea of “introducing routing in the cascading process” Dekoninck et al. (2024) and train an MDP for decision-making. Aggarwal et al. Aggarwal et al. (2023) directly determine both “whether to terminate the cascade and which stronger model to route to” within the MDP, whereas Hu et al. Hu et al. (2024a) only decide “which stronger model to route to” in the MDP and additionally rely on an answer-consistency strategy for cascading decisions.
4 Benchmarks and Applications
In this section, we briefly discuss popular benchmarks and applications in LLM Ensemble. First, there are two classic benchmarks specifically tailored for LLM Ensemble evaluation: 1) Benchmark MixInstruct proposed by Jiang et al. Jiang et al. (2023), which serves to assess the performance of LLM Ensemble, covers various instruction-following tasks and includes 11 popular large language models alongside 11K test samples; 2) Benchmark RouterBench proposed by Hu et al. Hu et al. (2024b), which can be used to evaluate methods that consider both performance and inference costs, especially suitable for ensemble-before-inference methods. Moreover, in terms of applications, beyond the methods outlined in Section 3, the concept of LLM Ensemble has found applications in a variety of more specialized tasks and domains. For instance, Lee et al.  Lee et al. (2023) leverage the ROUGE-L metric to assess the similarity of generated text and employed a similarity-based selection strategy for ensemble to produce Instruction-Tuning data; additionally, several studies focus on win rate evaluation Gao et al. (2024), SQL generation Gundabathula and Kolar (2024), and so on.
5 Discussion
5.1 Summarization
Here we provide a summary analysis of three LLM Ensemble paradigms and seven LLM Ensemble approaches, as illustrated in Table 4, considering the most critical methodological attributes: ensemble strategy, granularity and ensemble goal.
From the perspective of ensemble strategy, aggregation methods (involving the average or weighted ensemble of all model outputs) are more sophisticated compared to regeneration-based methods, which select a single output (equivalent to a hard voting process). However, regeneration-based methods are burdened by the additional need for large model-specific training data preparation and model training. From the perspective of ensemble granularity, it is apparent that response-level ensemble methods are relatively coarse-grained; while methods with finer granularity, particularly token-level ensemble methods, can more effectively harness the distribution information from each model during the decoding phase. Further, from the perspective of ensemble goals, (b) ensemble-during-inference methods and (c1) non-cascaded-based methods, unconstrained by cost considerations, can employ more flexible ensemble strategies beyond selection and utilize finer-grained ensemble approaches, ultimately yielding greater performance improvement potential.
5.2 Existing Limitations and Future Directions
Principled span-level ensemble-during-inference approach.
Ensemble-during-inference methods based on span-level ensemble already provide an adequate granularity. Nevertheless, current segment segmentation techniques are still overly simplistic (e.g., rigidly defining segment lengths as fixed values, such as a 4-word sequence). Shifting to a more principled segment segmentation approach could provide richer, more valuable information for the subsequent ensemble process.
Sophisticated unsupervised non-cascade ensemble-after-inference approach.
To achieve optimal performance, and in cases where frequent calls to the output distributions of different models during the decoding phase are not feasible, non-cascade ensemble-after-inference approach, which integrates the responses from multiple large models to derive the final answer, is highly practical. However, these methods are often limited either by relying on a simplistic pairwise similarity measure between different outputs, or requiring additional generation-oriented supervised learning components, thereby sacrificing generalization. Hence, the development of more sophisticated, unsupervised non-cascade ensemble-after-inference methods holds considerable importance.
General cascade approach.
Existing cascade methods, compared to ensemble-before-inference methods, can also address the cost-considered ensemble problem, with the added benefit of utilizing model responses during the cascade process to select the most appropriate output. However, most of these methods are not designed for generation tasks, and the only generation-oriented one relies on supervised learning, losing generalization. Thus, developing a general-purpose unsupervised cascade approach would mark significant progress.
6 Conclusion
LLM Ensemble stands as a direct manifestation of ensemble learning in the era of large language models. The accessibility, diversity, and out-of-the-box usability of large language models have made the spirit of ensemble learning shine even brighter, fueled by the rapidly developing LLM Ensemble studies. This paper provides a comprehensive taxonomy and a review of the methods in the field of LLM Ensemble, introduces relevant applications and benchmarks, conducts a summative analysis of these methods, and proposes several potential research directions. We hope that this review will offer valuable insights to researchers and inspire further advancements in LLM Ensemble and related research areas.